# Site-Specific Radio Channel Representation - Current State and Future Applications

### v2

@@ Abstract
A site-specific radio channel representation (SSCR) considers the surroundings of the communication system through the environment geometry, such as buildings, vegetation, and mobile objects including their material and surface properties. In this article, we focus on communication technologies for 5G and beyond that are increasingly able to exploit the specific environment geometry for both communication and sensing. We present methods for a SSCR that is spatially consistent, such that mobile transmitters (TXs) and receivers (RXs) cause a correlated time-varying channel impulse response. When modelled as random, this channel impulse response has non-stationary statistical properties, i.e., a time-variant Doppler spectrum, power delay profile, K-factor and spatial correlation. A SSCR will enable research into emerging 5G and beyond technologies such as distributed multiple-input multiple-output systems, reconfigurable intelligent surfaces, multi-band communication, and joint communication and sensing. These 5G and beyond technologies will be deployed for a wide range of environments, from dense urban areas to railways, road transportation, industrial automation, and unmanned aerial vehicles.
@@ I Introduction
An SSCR allows the accurate prediction of radio propagation to analyse the performance of communication and sensing systems for sites of general interest and for different vertical domains such as railways, road transport, manned and unmanned aircraft or robotic collaboration. In this paper, we are interested in reviewing models that can influence research, system design, and standardisation in different domains.
@@
Radio waves are emitted by the transmitting antenna and propagate via different propagation paths to the receiving antenna where they sum up linearly, causing dispersion in delay and Doppler. The resulting radio channel can be fully described by a sampled time-varying impulse response. The impulse response depends on the TX and RX coordinates, their velocity vectors as well as all objects in the environment. Its effect on the transmitted signal is computed by a time-varying convolution. Fundamentally, Maxwells equations can be solved numerically to obtain the impulse response, however the numerical complexity is too high for most practical cases. Hence, SSCRs aim to achieve a compromise between accuracy and complexity.
II Motivation for future standards and use cases
In the fall of 2022, the ComSoc Standards Development Board (DSB) initiated a new group on wireless propagation channel modeling, which in 2023 was formalized as the P1944 Standard under the aegis of the Mobile Communication Networks Standards Committee, with the Project Authorization Request approved by IEEE Nescom in May 2023. The special interest group is currently facilitating the development of a number of SSCRs for 5G and beyond described in the following subsections.
@@ II Motivation for future standards and use cases
It is important to distinguish the aims of P1944 from the channel modeling of 3GPP, 802.11/WiFi, and similar standards efforts. The latter aim to develop channel models best suited for the specific operating parameters and use cases of this standard. In contrast, P1944 aims to develop more generic models suitable for both academic research and industry developments, which might be used by standards organizations as basis for versions that are specialized to their purposes.
II-A Joint communication and sensing
@@
Sensing capabilities will create a disruptive and explosive growth of new 5G and beyond applications and services. These can be economically implemented at minimal or no additional cost in existing base stations and/or user equipment heralding a new 5G and beyond era of joint communications and sensing (JCAS) also known as integrated sensing and communication. This sensing capability will span a very wide range of applications as described in [1], including but not limited to: Intrusion detection, health monitoring, environmental monitoring, road traffic, daily life, industrial automation, etc. For example, a base station detects and alerts a pedestrian entering a pedestrian crossing while a car is about to enter, as shown in Fig. 1.
@@
Monostatic and bistatic network topologies are the simplest configuration for sensing. In monostatic sensing, a single base station implements the TX and RX functions at the same site, performing the sensing function within its radio coverage. In the bistatic case, two sites work together to achieve enhanced performance. The communication waveform may also be used opportunistically for the sensing function. The reliable and robust design of such solutions will require new propagation models that enable deployment in a variety of environments for different applications with a high probability of user satisfaction. Current communication models are excellent in the context of a TX-RX link. However, at a minimum, the return power ratios from the object being sensed and its surrounding background are critical enhancements to be added to the existing propagation modes. Other improvements required include spatial consistency and Doppler statistics.
II-B Railways
@@
Railway communications need to evolve from 2G (GSM rail, GSM-R) to the next generation railway communication system with improved performance and intelligence. 5G and related technologies [2] are seen as a solution to support the increasing data traffic, various new services, and high safety requirements for future intelligent railways. The next generation railway communication system is expected to support speeds of up to 500 km/h, which poses significant challenges for dynamic channel modelling, including dynamic estimation, tracking, and characterisation of multi-path components, which are not well studied for 2G railway communication systems and need to be addressed in the future. Channel dispersions in the delay, Doppler, and angle domain should also be accurately characterised by a SSCR for future intelligent railways for a wide range of scenarios such as viaducts, tunnels, cuttings, stations, etc. Well-designed large and small scale models are essential for site-specific coverage prediction and link budget calculation.
II-C Street transport
@@
Trains and vehicles are both examples of wireless channels where one or both end nodes are moving, so the channel models developed for these two use cases have many similarities, e.g., Doppler shift plays a major role and scattering objects on both sides of the track or road, such as buildings, foliage, uneven terrain, etc. must be considered. However, the movement of trains along a fixed track and the availability of utility poles at regular intervals make the case for trains generally more deterministic. In contrast, vehicle-to-everything (V2X) communication breaks down into many complex scenarios [3], ranging from static intra-vehicle or indoor parking to dynamic outdoor links between two vehicles or between a vehicle and the infrastructure. Vehicle radio channels are inherently location-specific. As an indicative example, the coherence time for a vehicular channel varies with the speed of the vehicle, the wavelength and the multipath distribution. There are more unique scenarios, such as an infrastructure-to-infrastructure channel where both TX and RX may be static, but moving vehicles acting as scatterers cause channel variations.
II-D Aircraft
@@
As the number of manned and unmanned aircraft continues to grow, there is a need to manage airspace more efficiently. This includes all aspects of communication, navigation and surveillance technologies. For all three aspects, it is important to understand the propagation channel between aircraft (manned or unmanned) and ground infrastructure. There are several key differences between the air-ground channel and traditional terrestrial communication systems. Aircraft can travel faster than terrestrial vehicles, both within and over areas where terrestrial communications take place. This poses challenges for accurate modelling of the air-ground channel, as Doppler spectra [4] and power delay profiles can be rapidly time-varying. Thus, time-varying Doppler and delay profiles provide non-stationary channel statistics.
@@
Site-specific air-ground channel conditions can manifest themselves when aircraft enter environments for which little modelling has been done. Examples include small aircraft conducting fly-by inspections of utilities such as large wind farms, advanced air mobility aircraft manoeuvring near rooftops in urban areas, and air-ground communications with a large swarm of aircraft. In addition, aircraft-specific effects such as airframe shadowing are not addressed in typical terrestrial channel modeling.
@@
In the common case where aircraft antennas are omnidirectional, the primary components of the channel impulse response are the line of sight and the earth surface reflection. The strength of the received line-of-sight and surface reflection also depends on the altitude of the aircraft, its bank angle or pitch during take-off and landing These components must be supplemented by additional multipath components (MPCs), which may be intermittent MPCs as the aircraft moves through a volume of space. In environments where the geometric and electrical properties of obstacles (or interacting objects) can be accurately quantified, these intermittent MPCs can be estimated and tracked in the channel model. In the case of propeller aircraft, the radio signal is additionally Doppler-shifted by the frequency of the revolutions of the rotating propeller.
II-E Large buildings and their construction
Given that of wireless traffic data is generated indoors, there is a strong need to model building entry loss, indoor obstruction by building structures, furniture and people. As an optional feature, indoor channel models can include the enhancements possible with strategically placed reconfigurable intelligent surfaces (RISs). A SSCRs will also play an important role in construction site automation with wireless technologies, taking into account the constant changes in building geometry during the construction phase, by an update of the geometry database, and subsequently in the operational phase of large infrastructures.
III Model type
@@
where is the time index sampled at time , is the frequency index and is the even number of samples in the frequency domain, denotes the direction of arrival vector in terms of azimuth and elevation of MPC at the TX side and the corresponding direction of departure at the RX side, respectively. The combined band-limited impulse response of TX and RX is denoted by . The normalised path delay is denoted by , the path weight by , and is the delay of path at time index .
The channel frequency response at the footpoint of the antennas, is finally obtained by integrating over the unit sphere surfaces and at TX and RX side, taking the antenna pattern of TX, , and RX, , into account
@@
SSCRs consider non-stationary radio propagation conditions. In all practical scenarios the radio channel exhibits a local wide-sense stationarity for a region in time and frequency [3]. This property allows to simplify the SSCR such that the MPC parameters are constant for a stationarity region. The computation of the MPC parameters can be done in several site-specific ways, as described below.
III-A Ray tracing
Ray tracing is a high-frequency approximation used to solve Maxwell’s equations. It has long been used successfully for deterministic channel prediction [6], particularly for deployment planning, both indoor and outdoor. As it computes the parameters of the MPCs, including amplitude, direction of arrival, and direction of departure, it provides a true double-directional channel characterization that can be combined with (almost) arbitrary antenna patterns and array patterns in post-processing. In addition, ray tracing results are inherently spatially consistent, i.e., the change in power, angles, and delays follows the environment’s geometry, and - in a static environment - the same user equipment location is associated with the same channel. The requirements of 5G and beyond have stimulated research in four directions: distributed multiple-input multiple-output (D-MIMO), ultrahigh frequencies, and RISs.
@@ III-A Ray tracing
Ways to increase efficiency include:
@@
III-B Quasi-deterministic channel model
Quasi-deterministic models combine deterministic predictions, e.g., from a ray tracer, with stochastic channel modeling. In one incarnation, simplified ray tracing (a small number of reflections considered) is performed on a low-resolution environmental map to deterministically find the dominant propagation paths. Smaller MPCs associated with the dominant contributions are then generated stochastically (but with their delay/angle determined by the dominant paths). The delay/angle positions and amplitudes of additional clusters of MPCs can be generated purely stochastically to (i) represent static objects such as trees, (ii) represent moving objects such as pedestrians and cars, or (iii) model diffuse reflection on rough surfaces; in the second case, the birth/death statistics of the associated MPCs are also generated stochastically. In a related approach, the cluster centers are determined from the (rough) geometry of the environment. The associated intra-cluster statistics, usually obtained from measurements, are then provided for a complete channel description. Quasi-deterministic models [9] were first proposed in the early 2000s, and their concepts were then re-used by the European METIS project and adopted as an option for 3GPP simulations.
@@ III-B Quasi-deterministic channel model
Compared to purely stochastic models (such as 3GPP), quasi-deterministic models have the advantages of spatial consistency (since the evolution of angles and delays of clusters or dominant paths follows from the geometry) and easier incorporation of specific geometries (e.g., the impact of street width can be easily incorporated). The advantages compared to ray tracing are the significantly reduced computational effort, easier exchange of environmental databases between entities that want to compare their results, and a straightforward way of incorporating mobile scatterers.
Figure 3 shows an exemplary quasi-deterministic environment model for a vehicular scenario including buildings, vegetation, parked cars, and city lamp posts. The geometry model contains open street map data augmented by objects detected via LIDAR and video data. It was validated with empirical radio channel measurements and frame error rate measurements.
III-C AI/ML based channel modeling

### v4

# Site-Specific Radio Channel Representation for 5G and 6G

@@ Abstract
A site-specific radio channel representation (SSCR) takes the surroundings of the communication system into account by considering the environment geometry, including buildings, vegetation, and mobile objects with their material and surface properties. We present methods for an SSCR that is spatially consistent, such that mobile transmitter and receiver cause a correlated time-varying channel impulse response and closely spaced antennas are correctly correlated. An SSCR is composed of a dynamically varying number of multipath components solely defined by the environment geometry and the material of the environmental objects. Hence, the environment geometry is the only natural scenario parameterization and specific calibration procedures shall be avoided. 5G and 6G physical layer technologies are increasingly able to exploit the properties of a wide range of environments from dense urban areas to railways, road transportation, industrial automation, and unmanned aerial vehicles. The channel impulse response in this wide range of scenarios has generally non-stationary statistical properties, i.e., the Doppler spectrum, power delay profile, K-factor and spatial correlation are all spatially variant (or time-variant for mobile receivers). SSCRs will enable research and development of emerging 5G and 6G technologies such as distributed multiple-input multiple-output systems, reconfigurable intelligent surfaces, multi-band communication, and joint communication and sensing. We highlight the state of the art and summarize research directions for future work towards an SSCR.
@@ I Introduction
A site-specific radio channel representation (SSCR) allows the accurate prediction of radio propagation to analyze the performance of communication and sensing systems for sites of general interest and for different vertical domains such as railways, road transport, manned and unmanned aircraft or robotic collaboration. In this paper, we are interested in reviewing models that can influence research, system design, and standardization in different domains.
@@
Radio waves are emitted by the transmitting antenna and propagate via different propagation paths to the receiving antenna where they sum up, causing dispersion in delay and Doppler. The resulting radio channel can be fully described by a sampled double directional time-varying impulse response. The impulse response depends on the transmitter (TX) and receiver (RX) coordinates, their velocity vectors as well as all objects in the environment. Its effect on the transmitted signal is computed by a time-varying convolution.
Maxwell’s equations can be solved numerically to obtain the impulse response; however, the numerical complexity is too high for most practical cases. Therefore, an SSCR must identify a balance between precision and computational burden. The accuracy of an SSCR improves with a more comprehensive geometry database and the incorporation of additional propagation effects, such as the reflection order, diffuse surface reflection, and diffraction effects. This paper will elucidate the currently known methods and algorithms to reduce the computational complexity of SSCR, achieving even real-time operation, while still aligning with empirical measurement data.
Organization of the paper
In Section II we will motivate the need for an SSCR that is able to cover research and development work for upcoming new physical layer technologies as described in Sec. III and new application scenarios as explained in Sec. IV. Three basic methods for an SSCR are introduced in Sec. V, which use the scenario geometry as natural parameterization and avoids additional calibration steps. In Sec. VI we summarize a step-by-step description for the setup, validation, and use of an SSCR. Open research directions for model validation techniques are introduced in Sec. VII and unexplored channel measurement scenarios are outlined in Sec. VIII. In Sec. IX real-time radio channel emulation method for an SSCR are explained. Final conclusions are drawn and open research directions are summarized in Sec. X.
II Motivation for a site-specific channel model
The current channel model for the new radio (NR) physical layer of 5G and 6G is defined in [1]. It provides methods for link and system level simulations focusing on a small set of scenarios such as urban microcell, street canyon, urban macrocell, indoor office, rural macrocell, and indoor factory. The methods defined by 3GPP in [1] have a long history starting with GSM in the 1990s where a narrowband voice transmission channel could be well modeled using a power delay profile and a Doppler spectrum per channel tap.
Over the last thirty years the channel modeling framework has been expanded to handle new physical layer features from 3G to 5G, such as wideband transmission, multiple-input multiple-output (MIMO) techniques, and new frequency bands. The current version of [1] allows for geometry based modeling but considers only the first and last interaction point and the delay of a multipath component is not defined by the geometry. To achieve spatial and frequency consistency, specific algorithms and parameters are provided for each scenario. Clearly [1] aims at computational efficiency enabling link and system level result comparisons.
With upcoming new physical layer technologies such as joint communication and sensing (JCAS), distributed multiple-input multiple-output (D-MIMO) systems and reconfigurable intelligent surfaces (RISs) a more comprehensive SSCR is required, that (a) includes scenarios with non-stationary propagation properties such as vehicles on railways, streets, and in the air and (b) also allows for a wide aperture of antenna system (i.e., the near-field extends to multiple ).
Hence, an SSCR aims at a fundamental numerical modeling approach that is geometry based and can leverage the computing power available on modern graphical processing uni platforms as well as data driven methods. An SSCR uses a dynamically varying number of multipath components (MPCs) solely defined by the environment geometry and the materials of the objects in the environment. An SSCR shall inherently provide spatial and frequency consistent channel impulse responses, taking the geometry and the material properties of the environment objects into account without the definition of special procedures and parameters for each scenario.
@@ II Motivation for a site-specific channel model
The long term work on an SSCR is performed in the P1944 Standard working group under the aegis of the Mobile Communication Networks Standards Committee. It is important to distinguish the aims of P1944 from the channel modeling of 3GPP, 802.11/WiFi, and similar standards efforts. The latter aim to develop channel models best suited for the specific operating parameters and use cases of a standard. In contrast, P1944 aims to develop more generic models suitable for both academic research and industry developments, which might be used by standards organizations as basis for versions that are specialized to their purposes.
III New physical layer technologies
Channel estimation is a fundamental challenge for any coherent communication system. The channel estimation algorithm needs to work reliably in all propagation scenarios. For the research and development of channel estimation methods, a link-level simulation with a realistic radio channel model is a key pre-requisite. For 5G and 6G technologies an SSCR becomes a fundamental requirement to enable research and development in the areas described below.
III-A Joint communication and sensing
@@
Sensing capabilities will create a disruptive and explosive growth of new 5G and 6G applications and services. These can be economically implemented at minimal or no additional cost in existing base stations and/or user equipment heralding a new 5G and 6G era of JCAS (also known as integrated sensing and communication). This sensing capability will span a very wide range of applications, including but not limited to: Intrusion detection, health monitoring, environmental monitoring, road traffic, daily life, industrial automation, etc. For example, a base station detects and alerts a pedestrian entering a pedestrian crossing while a car is about to enter, as shown in Fig. 1.
@@
Mono-static and bi-static network topologies are the simplest configuration for sensing. In mono-static sensing, a single base station implements the TX and RX functions at the same site, performing the sensing function within its radio coverage. In the bi-static case, two sites work together to achieve enhanced performance. The communication waveform may also be used opportunistically for the sensing function. The reliable and robust design of such solutions will require new propagation models that enable deployment in a variety of environments for different applications with a high probability of user satisfaction. Current radio channel models are excellent in the context of a TX-RX link. However, at a minimum, the return power ratios from the object being sensed and its surrounding background are critical enhancements to be added to the existing propagation modes. Other improvements required are to include spatial consistency and Doppler statistics.
III-B Antenna element group with wide aperture
With emerging technologies such as RIS and D-MIMO, antenna elements are distributed over a large area. For RISs potentially the whole facade of a building can be covered by spaced antenna elements with adjustable reflection coefficients. For D-MIMO, distributed remote radio heads can be deployed over an area with the size of about one , being connected to a central processing unit (preferably with fibre optic links) that processes their signals jointly and coherently. In both cases an SSCR is required that models the near field characterized by spherical wave fronts.
III-C Multi-band communication
With decreasing wave-length of the TX the received power decreases quadratically. Hence, directive antennas are required for frequency range (FR) 2 ( GHz to GHz) and FR3 ( GHz to GHz) in 5G and 6G systems. The required beam search to establish a stable link between TX and RX is a challenging and time-consuming task in outdoor mobile scenarios. Hence, multi-band communication aims to use a radio signal, e.g. in FR1 ( to GHz), to identify a suitable direction of arrival and direction of departure. These directions are used to establish a high-data rate broadband link in FR2 or FR3. For this technology a frequency-consistent SSCR is needed.
IV Application scenarios
For the reliable communication with vehicles to enable remote control and safety applications, the wireless communication system must be able to handle non-stationary propagation conditions for the drive-by case. Such non-stationary transitions are not available in classic models such as [1] and require an SSCR. Below we explore these future use cases in more detail and highlight scenario specific aspect that needs to be reflected by an SSCR.
IV-A Railways
@@
Railway communications must evolve from 2G (GSM rail, GSM-R) to the next generation railway communication system with improved performance and intelligence. 5G and related technologies [2] are seen as a solution to support the increasing data traffic, various new services, and high safety requirements for future intelligent railways. The next generation railway communication system is expected to support speeds of up to 500 km/h, which poses significant challenges for dynamic channel modeling, including dynamic estimation, tracking, and characterization of MPCs, which are not well studied for 2G railway communication systems and need to be addressed in the future. Channel dispersion in the delay, Doppler, and angle domain should also be accurately characterized by an SSCR for future intelligent railways for a wide range of scenarios such as viaducts, tunnels, cuttings, stations, etc. For characteristic environments encountered in railway scenarios, an SSCR demonstrates distinct responses tailored to each specific scene. For instance, in constrained environment like tunnels, it needs to account for the confined space and multiple reflections. In contrast, in open areas like viaducts, the focus might be on line-of-sight and diffraction over edges. Well-designed large and small scale models are essential for site-specific coverage prediction and link budget calculation. By tailoring specific methods and parameters to specific environments, an SSCR can provide highly accurate and realistic simulations of radio wave propagation, enabling better design and optimization of wireless communication systems.
IV-B Street transport
@@
Vehicular communication on streets is defined by a wireless channel where potentially both, TX and RX, are moving. Hence, time-variant Doppler shifts play a major role for an SSCR and scattering objects on both sides of the road, such as buildings, foliage, uneven terrain, etc. must be considered. Vehicular communication breaks down into different scenarios [3], ranging from static intra-vehicle or indoor parking to dynamic outdoor links between two vehicles or between a vehicle and the infrastructure. Vehicular radio channels are inherently location-specific resulting in a non-stationary fading process defined by the vehicle movement, and the movement of the surrounding vehicles in closer proximity. The building structure, street width and vegetation along the street have a strong influence on the radio channel properties as summarized in [3].
IV-C Aircraft
@@
As the number of manned and unmanned aircraft continues to grow, there is a need to manage airspace more efficiently. This includes all aspects of communication, navigation and surveillance technologies. For all three aspects, it is important to understand the propagation channel between aircraft (manned or unmanned) and ground infrastructure. There are several key differences between the air-ground channel and traditional terrestrial communication systems. Aircraft can travel faster than terrestrial vehicles, both within and over areas where terrestrial communications take place. This poses challenges for accurate modeling of the air-ground channel, as the Doppler spectral density (DSD) [4] and power delay profile (PDP) can be rapidly time-varying.
@@
Site-specific air-ground channel models are required, e.g., for small aircraft conducting fly-by inspections of utilities such as large wind farms, advanced air mobility aircraft maneuvering near rooftops in urban areas, and air-ground communications with a large swarm of aircraft. In addition, aircraft-specific effects such as airframe shadowing are not addressed in typical terrestrial channel modeling.
@@
In the common case where aircraft antennas are omnidirectional, the primary components of the channel impulse response are the line of sight and the earth surface reflection. The strength of the received line-of-sight and surface reflection also depends on the altitude of the aircraft, its bank angle or pitch during take-off and landing. These components must be supplemented by additional MPCs, which may be intermittent MPCs as the aircraft moves through a volume of space. In environments where the geometric and electrical properties of obstacles (or interacting objects) can be accurately quantified, these intermittent MPCs can be estimated and tracked in the channel model. In the case of propeller aircraft, the radio signal is additionally Doppler-shifted by the frequency of the revolutions of the rotating propeller.
V Methods for site specific channel representation
@@
where is the time index sampled at time , is the frequency index and is the even number of samples in the frequency domain, denotes the direction of arrival vector in terms of azimuth and elevation of MPC at the TX side and the corresponding direction of departure at the RX side, respectively. The combined band-limited impulse response of the TX and RX hardware is denoted by . The normalized path delay is denoted by , the path weight by , and is the delay of path at time index .
The channel frequency response at the base point of the antennas, is finally obtained by integrating over the unit sphere surfaces and at TX and RX side, taking the antenna pattern of TX, , and RX, , into account
@@
SSCRs consider non-stationary radio propagation conditions. In all practical scenarios the radio channel exhibits a local wide-sense stationarity for a region in time and frequency [3]. This property allows to simplify the SSCR such that the MPC parameters are constant for a stationarity region. The computation of the MPC parameters can be done with different site-specific methods, as described below. These methods aim on general radio channel representation avoiding scenario specific parameterization or calibration procedures.
V-A Ray tracing
Ray tracing is a high-frequency approximation used to solve Maxwell’s equations. It has long been used successfully for deterministic channel prediction [6], particularly for deployment planning, both indoor and outdoor. As it computes the parameters of the MPCs, including amplitude, direction of arrival, and direction of departure, it provides a true double-directional channel characterization that can be combined with (almost) arbitrary antenna patterns and array patterns in post-processing. In addition, ray tracing results are inherently spatially consistent, i.e., the change in power, angles, and delays follows the environment’s geometry, and - in a static environment - the same user equipment location is associated with the same channel.
@@ V-A Ray tracing
Open research areas for ways to increase efficiency include:
@@
V-B Quasi-deterministic method
Quasi-deterministic SSCRs, also called geometry based stochastic channel models (GSCMs) in the literature, combine deterministic predictions, e.g., from a ray tracer, with stochastic but still geometry based channel modeling.
Simplified ray tracing (a small number of reflections considered) is performed on a low-resolution environmental map to deterministically find the dominant MPC (the map can be actual or synthetically generated). Smaller MPCs associated with the dominant contributions are then generated stochastically either (i) in the geometric space by placing point scatterers according to a given distribution close to surfaces of geometric objects to represent surface roughness and point clusters to represent other objects such as vegetation, moving vehicles, and pedestrian or (ii) in the delay and direction of departure/direction of arrival domain. In both cases birth/death statistics of the associated MPCs and intra-cluster statistics are usually obtained from measurements.
@@ V-B Quasi-deterministic method
Quasi-deterministic models were first proposed by Kunisch and Pamp and by COST 259 in the early 2000s, and their concepts were then re-used by the European METIS and MiWeBa projects and adopted as an option for 3GPP simulations [9]. Compared to purely stochastic models (such as 3GPP), quasi-deterministic models have the advantages of spatial consistency (since the evolution of angles and delays of clusters or dominant paths follows from the geometry) and easier incorporation of specific geometries (e.g., the impact of street width can be easily incorporated). The advantages compared to ray tracing are the significantly reduced computational effort, easier exchange of environmental databases between entities that want to compare their results, and a straightforward way of incorporating mobile scatterers.
V-C AI/ML based method
A promising approach of SSCR is data-driven AI/ML-based channel modeling [10]. AI/ML-based channel models are trained to learn the mapping relationship between complex environmental features and channel characteristics. It can achieve a balance between generalization and prediction accuracy compared to traditional channel models. An SSCR requires the prediction of channel characteristics of a specific site according to environmental conditions, and for AI/ML-based modeling, the ability to extract features from environmental information determines the prediction accuracy of the model.
For the extraction of environmental features, different data sources can be used, such as images, satellite maps, point cloud data, etc. The impact of data resolution on the accuracy of environmental features should be carefully considered. Methods like neural networks are usually used to extract propagation-related features from the data, which are further employed to predict channel characteristics. In the phase of model selection and network training, supervised learning is the most common approach, where models are trained to minimize prediction errors using labeled data, unsupervised learning methods are also employed in certain scenarios. For example, clustering algorithms can classify different channel conditions, and auto-encoders can extract meaningful features from raw data without labels. Additionally, generative adversarial networks can generate synthetic channel data, which is useful in scenarios with limited real-world data.